OpenStudy (anonymous):

correct me if i did this wrong 6x+2x-8+5=7-2x 8x-3=7-2x 8x+2x=7+13 10x=20 x=2

OpenStudy (abb0t):

Looks right to me.

OpenStudy (anonymous):

Instead of adding 13 to both sides, you should have added 3 in the third step. This would leave you with 10x=10 in which x=1

OpenStudy (anonymous):

so did i do it wrong @escofield784

OpenStudy (anonymous):

and can you explain this one im just practicing, like i did this one i showed you but im stuck on this -2x - 18 = 6(1 - 2x) + 4x x =

OpenStudy (anonymous):

distribute the 6 first. This will give you -2x-18=6-12x+4x -2x-18=6-8x 6x-18=6 Add 8x to both sides 6x=24 Add 18 to both sides x=4 Divide both sides by 6

OpenStudy (anonymous):

thanks i was doing it complety different i added 2x to both sides